OpenStudy (anonymous):

katrina paid $25. for a box fan that was marked down 15%. what was the original price

OpenStudy (anonymous):

let original price = x x - 15% of x = 25 ==> x - 0.15x = 25 ==> 0.85x = 25 ==> x = 25/0.85 ==> x = 29.41 approx

OpenStudy (anonymous):

how did you get the 0.85x = 25? that's where I always get stuck

OpenStudy (anonymous):

x means 1x so 1x - 0.15x = 0.85x

OpenStudy (anonymous):

i appreciate your help :)